# HMIL: Hierarchical Multi-Instance Learning for Fine-Grained Whole Slide Image Classification

### v1

@@ 2.1 MIL in WSI Classification
In WSI classification, to tackle the challenges of gigapixel-sized WSIs with weak annotations, many methods primarily the MIL framework. This framework involves three main stages: extracting features at the patch level, aggregating these patch-level features into slide-level representations, and training a classifier with these representations using slide labels for fully supervised prediction. Existing MIL methods in WSI classification can be broadly categorized based on their reliance on instance-level annotations.
@@ 4.3.1 Fine-grained Classification
Figure 4: The class-wise AUC distribution of top-performing methods on BRACS (Left), PANDA (Middle), and CCC (Right) datasets.
@@ 5 Discussion and Conclusion
Our experiments underscore the mutual benefits of integrating hierarchical alignment in bolstering both fine-grained and coarse-grained WSI classification. In the current study, we introduced hierarchical guidance through class-wise attention learning, demonstrating its potential. However, it must be acknowledged that fine-grained classification still has a long way to go, particularly because the sensitivity is not yet sufficiently high. Recent state-space model-based approaches have shown promise in multiple-instance learning (MIL) applications due to their long-range sequence modeling capabilities, achieving competitive results in our studies. Future work could explore the incorporation of hierarchical guidance into these MIL architectures [59, 40, 60]. Additionally, we should not neglect the domain bias present in WSIs, which can arise from inference factors such as background noise, z-stack artifacts, and tissue or cell distortion. Investigating domain adaptation and debiasing techniques could enhance model robustness and generalizability, particularly when dealing with diverse and heterogeneous imaging modalities [61]. Overall, these insights highlight the importance of both hierarchical guidance and addressing domain biases in advancing fine-grained WSI classification methods.
In summary, we introduced the HMIL framework, an approach that leverages label hierarchy into the MIL framework to address the fine-grained WSI classification task. HMIL comprehensively aligns features from the instance to the bag level. The framework further incorporates dynamic weighting and supervised contrastive learning, which refine slide-level representations, resulting in improved classification outcomes. Crucially, HMIL eliminates the need for extensive instance-level annotations. It demonstrates robust performance across various WSI datasets from different imaging modalities, including the publicly available histology datasets BRACS and PANDA, as well as our extensive cervical cytology WSI dataset CCC, underscoring its effectiveness and adaptability.

### v3

@@ 2.1 MIL in WSI Classification
In WSI classification, to tackle the challenges of gigapixel-sized WSIs with weak annotations, many methods primarily utilize the MIL framework. This framework involves three main stages: extracting features at the patch level, aggregating these patch-level features into slide-level representations, and training a classifier with these representations using slide labels for fully supervised prediction. Existing MIL methods in WSI classification can be broadly categorized based on their reliance on instance-level annotations.
@@ 4.3.1 Fine-grained Classification
Figure 4: The class-wise AUC distribution of top-performing methods on BRACS (Top), PANDA (Middle), and CCC (Bottom) datasets.
@@ 5 Discussion and Conclusion
Our experiments demonstrate the efficacy of hierarchical alignment in enhancing both fine-grained and coarse-grained WSI classification through class-wise attention learning. However, the limitations in fine-grained classification sensitivity necessitate further investigation. At the feature extraction stage, recent pathology foundation models which pretrained on extensive pathological datasets [59, 60, 61], offer potential for more representative feature extraction and semantic interpretation. Integrating these models into our hierarchical framework or implementing hierarchical pretraining techniques [62] may potentially improve the current classification performance. At feature aggregation stage, state-space model-based approaches have exhibited encouraging results in MIL applications through their superior sequence modeling capabilities. Future research should explore incorporating hierarchical guidance into these architectures [63, 40, 64]. Moreover, addressing domain bias arising from imaging artifacts and tissue variations through domain adaptation and debiasing techniques is crucial for model robustness and generalizability across heterogeneous imaging modalities [65]. These insights underscore the synergistic importance of hierarchical guidance and domain adaptation in advancing WSI classification methodology.
In this work, we introduced the HMIL framework, an approach that leverages label hierarchy into the MIL framework to address the fine-grained WSI classification task. HMIL comprehensively aligns features from the instance to the bag level. The framework further incorporates dynamic weighting and supervised contrastive learning, which refine slide-level representations, resulting in improved classification outcomes. Crucially, HMIL eliminates the need for extensive instance-level annotations. It demonstrates robust performance across various WSI datasets from different imaging modalities, including the publicly available histology datasets BRACS and PANDA, as well as our extensive cervical cytology WSI dataset CCC, underscoring its effectiveness and adaptability.